10 kg water falls through a distance of 854m and all the energy is effective in heating the water.To what temperature will the water be raised if it was initially at 20c?


1
Expert's answer
2021-03-14T19:21:53-0400

"mgh=cm\Delta t,"

"gh=c(t_2-t_1),"

"t_2=t_1+\frac{gh}{c},"

"t_2=20+\frac{9.8\cdot 854}{4200}=22°C."